Play ebruary Und data Production Sales Variable costs S 900 S 900 00 Foxed costsCosts S400,000 $00,000 $400,000 S140,000 $140,000 $140,000 costs budgeted level of production used to calculate the budgeted variances. variance is written off to cost of goods sold in the month in which it occurs for BigScreen in January, February, and March of 2012 under (a) variable 76%
Answer #1

January

February

March

Beginning inventory

0

300

300

Production

1000

800

1250

Goods available for sale

1000

1100

1550

Units sold

700

800

1500

Ending inventory

300

300

50

The budgeted fixed manufacturing cost per unit and budgeted total manufacturing cost per unit under absorption costing are:

January

February

March

(a)

Budgeted fixed manufacturing costs

400000

400000

400000

(b)

Budgeted production

1000

1000

1000

(c)=(a)÷(b)

Budgeted fixed manufacturing cost per unit

400

400

400

(d)

Budgeted variable manufacturing cost per unit

900

900

900

(e)=(c)+(d)

Budgeted total manufacturing cost per unit

1300

1300

1300

Part 1

Variable costing

January 2012

February 2012

March 2012

Revenues

1750000

2000000

3750000

Variable costs

Beginning inventory

0

270000

270000

Variable manufacturing costs

900000

720000

1125000

Cost of goods available for sale

900000

990000

1395000

Deduct ending inventory

(270000)

(270000)

(45000)

Variable cost of goods sold

630000

720000

1350000

Variable operating costs

420000

480000

900000

Total variable costs

1050000

1200000

2250000

Contribution margin

700000

800000

1500000

Fixed costs

Fixed manufacturing costs

400000

400000

400000

Fixed operating costs

140000

140000

140000

Total fixed costs

540000

540000

540000

Operating income

160000

260000

960000

2500*700 =1750000

2500*800 =2000000

2500*1500 =3750000

300*900 = 270000

1000*900= 900000

800*900= 720000

1250*900 = 1125000

300*900 = 270000

50*900 = 45000

600*700 = 420000

600*800 = 480000

600*1500 = 900000

B Absorption costing

January 2012

February 2012

March 2012

Revenues

1750000

2000000

3750000

Cost of goods sold

Beginning inventory

0

390000

390000

Variable manufacturing costs

900000

720000

1125000

Allocated fixed manufacturing costs

400000

320000

500000

Cost of goods available for sale

1300000

14300000

2015000

Deduct ending inventory

(390000)

(390000)

(65000)

Adjustment for prod. vol. var

0

80000

(225000)

Cost of goods sold

910000

1120000

1725000

Gross margin

840000

880000

2025000

Operating costs

Variable operating costs

420000

480000

900000

Fixed operating costs

140000

140000

140000

Total operating costs

560000

620000

1040000

Operating income

280000

260000

985000

300*1300 = 390000                   

1000*400 = 400000

800*400 = 320000

1250*400 = 500000

300*1300 = 390000

50*1300 = 65000

900000-720000 = 80000

900000-1125000 = -225000

Part 2

January = 280000-160000 = 120000

February = 260000-260000 = 0

March = 985000-960000 = 25000

Moving fixed manufacturing costs into inventories is the main reason for the difference between absorption and variable costing